A coil is connected in series with a 14.0 kΩ resistor. An ideal 56.3 V battery is applied across the two devices, and the current reaches a value of 2.20 mA after 5.59 ms. (a) Find the inductance of the coil. (b) How much energy is stored in the coil at this same moment?
Answer #1

b) energy stored

E = (1/2) L i​​​​​​2 = (1/2)×98.8×(2.2×10-3)2 = 2.39 × 10-4 J

= 0.239 mJ